# CoRe: Context-Robust Remasking for Diffusion Language Models

### v2

@@ Abstract
Keywords Diffusion language models $\cdot$ Inference-time revision $\cdot$ Robustness
@@ 1 Introduction
Masked Diffusion Models (MDMs) have emerged as a promising alternative to autoregressive decoding for discrete sequence modeling [28], allowing parallel token updates through iterative unmasking [25]. At inference time, an MDM starts with a fully masked sequence and progressively constructs a response. At each diffusion step, the sampler predicts token distributions for masked positions conditioned on the current partially unmasked sequence, selects a subset of positions to umask, and unmasks them. This iterative process means that unmasked tokens in the early steps are selected under a partial context that evolves as decoding proceeds. To improve quality, some variants employ revision strategies, remasking a small subset of previously unmasked tokens to allow resampling [35].
@@ 2 Related Work
Crucially, these revision methods either rely on unreliable proxy quality signals or require additional training. ReMDM’s standard strategy uses stale confidence scores from past decoding steps, ignoring updated context. P2-Self uses current token probabilities, which can be misleadingly high even for coherent but incorrect generations. Moreover, static confidence scores are often uncalibrated [14, 18]. Alternatively, other approaches, such as PRISM [15], RemeDi [13], or GIDD [vonrütte2025gidd], learn specialized policies to guide revision. Steering or reward-guided methods [27, 34, 24] optimize auxiliary objectives to guide sampling. While effective, these methods require training or incur significant overhead (e.g., ensembling or particle resampling [20, 30]), preventing plug-and-play use. In contrast, CoRe is context-aware and training-free. Rather than trusting static or stale token uncertainty metrics, we identify context-brittle tokens—those whose likelihood is not stable under context perturbations—framing revision as a distributionally robust optimization problem.
@@ 3 Problem Formulation
Let $\mathcal{V}$ be a vocabulary. Given a prompt $x$, the MDM decoder’s state at step $t$ is represented by $y^{(t)}=[x;\ y^{(t)}_{1},\ldots,y^{(t)}_{L}\,$], where each $y^{(t)}_{i}\!\in\!\mathcal{V}\cup\{\texttt{[MASK]}\}$. Decoding initializes the $L$ response positions $y^{(t)}_{1},\ldots,y^{(t)}_{L}$ to [MASK]and progressively replacing masks with tokens over steps $t=1,\ldots,T$. At each step, for every masked position $i$, the model computes:
$p_{\theta}(y_{i}^{(t)}=v\mid y^{(t)},i),\quad v\in\mathcal{V},$
and the sampler selects which positions (and how many) to unmask, producing $y^{(t+1)}$.
@@ 4 Method
Figure 1: Illustration of Context-Robust Remasking (CoRe). Our method operates on the current state $y^{(t)}$, where the response is partially unmasked. (Left) Selection: We select potentially brittle unmasked tokens (red dashed box) to test for stability, distinct from the next token scheduled for unmasking (blue dashed box). (Top-Right) CoRe Mechanism: We mask the selected tokens to create a Perturbed Context $\tilde{y}^{(t)}$ and then compute their Instability Scores under this new context. The token “a” is found to be the most brittle (high instability) and is updated to “an,” which is the most likely token given the perturbed context. (Bottom-Right) Base Unmasking: In parallel, the base model uses the original context $y^{(t)}$ to predict the next token (“icy”); this newly unmasked token (“icy”) is combined with the updated token (“an”) to form the Next State $y^{(t+1)}$, yielding the contextually consistent phrase “an icy.”
@@ 4.1 Context-Robust Token Remasking Framework
For a step $t$, let $C_{t}$ be the set of unmasked indices eligible for revision. Let $S_{t}\subseteq C_{t}$ be the indices selected for perturbation (using a selection strategy detailed in Sec. 4.2). We define the perturbed context $\tilde{y}^{(t)}$ by masking the positions in $S_{t}$:
@@
This operation occludes the information at indices $S_{t}$, forcing the model to rely on the remaining context. In our framework, we formally define these masking operations as perturbations to the information content, allowing us to cast revision as a robustness optimization problem.
@@
For each token position $i\in S_{t}$, we define the instability score as the negative log-likelihood of the current token $y^{(t)}_{i}$ under the perturbed context $\tilde{y}^{(t)}$:
$\ell_{i}\;\triangleq\;-\log p_{\theta}\!\left(y^{(t)}_{i}\mid\tilde{y}^{(t)}\right).$ (2)
A high instability score indicates that $y_{i}^{(t)}$ is context-brittle: although the token was originally generated with high confidence, it is unsupported by the perturbed context $\tilde{y}^{(t)}$. In practice, we use $\ell_{i}$ to rank candidate positions $i\in S_{t}$: positions with larger $\ell_{i}$ are more sensitive to context perturbation (masking $S_{t}$) and are prioritized for revision. The average instability score over the candidate positions is then:
@@
The masking operation in Eq. (1) produces a family of nearby contexts obtained by removing information from a small set of token positions, which varies with different $S_{t}$. To formalize the search over these contexts, we introduce a binary mask variable $z\in\{0,1\}^{L}$, where setting $z_{i}=1$ corresponds to masking token $i$; for indices $i\notin C_{t}$ we fix $z_{i}=0$ (non-revisable positions are never perturbed). We define the distribution over these masks as Bernoulli trials:
@@
where $\pi_{i}$ controls the probability of perturbing position $i$, and $m$ bounds the expected number of perturbed tokens. For any fixed token $a\in\mathcal{V}$, the Dirac delta distribution $\delta_{a}(\cdot)$ is defined as:
@@
This defines a probability distribution that assigns unit mass to the single token $a$. Sampling from $\delta_{a}$ always returns $a$ with probability one.
@@
Given a mask $z$, the perturbed sequence $\tilde{y}$ is generated independently across token positions as:
@@
If $z_{i}=0$, the factor $\delta_{y^{(t)}_{i}}(\tilde{y}_{i})$ ensures that the token remains unchanged, i.e., $\tilde{y}_{i}=y^{(t)}_{i}$. If $z_{i}=1$, the factor $\delta_{\texttt{[MASK]}}(\tilde{y}_{i})$ enforces the deterministic replacement of the token with the mask token. $S_{t}=\{i\mid z_{i}=1\}$ denotes the set of indices masked by $z$. Marginalization over $z$ yields a distribution over masked sequences $\tilde{y}$:
@@
Step 1: Worst-Case Context Perturbation. The objective for achieving the worst-case perturbed context distribution at decoding step $t$ is defined as the distribution that achieves the highest instability scores:
$\pi^{*}=\operatorname*{arg\,max}_{\pi\in\Pi}\;\Big[\mathbb{E}_{\tilde{y}\sim Q_{\pi}(\tilde{y})}\mathcal{L}(S_{t})\Big],\qquad\text{s.t.}\;\;\Pi=\Big\{\pi\in[0,1]^{|C_{t}|}:\sum_{i\in C_{t}}\pi_{i}\leq m\Big\},$ (8)
where $S_{t}$ denotes the set of indices masked in $\tilde{y}$.
[…]
@@
$y^{(t)}_{i}\leftarrow\operatorname*{arg\,min}_{v\in\mathcal{V}}\Big[-\log p_{\theta}\!\left(v\mid\tilde{y}^{(t)}\right)\Big],\quad\forall i\in\mathcal{I}_{t}.$ (10)
@@ 4.2 An Efficient Remasking Algorithm
Directly optimizing the worst-case context perturbation in Eq. (8) requires searching over all subsets of token indices and is combinatorially intractable. We therefore consider a single deterministic mask—i.e., a candidate index set $S_{t}$ to perturb. We show in Appendix C that the instability computed on this subset $S_{t}$ constitutes a valid lower bound on the worst-case instability objective. To make this choice tractable, we construct $S_{t}$ based on token contention, specifically selecting the positions with the smallest top-2 probability margins:
@@
where $v_{i}^{1}$ and $v_{i}^{2}$ denote the tokens with the highest and second-highest probabilities. Prior work indicates that small top-2 margins reliably signal token contention: competing values are close in probability [17]; in such cases, even slight context shifts can flip the top-1 prediction. We therefore use the margin to select the candidate subset $S_{t}$. Crucially, while margin identifies where to probe, the revision itself is determined solely by the instability score computed under the perturbed context. This perturbation serves a dual purpose: it quantifies brittleness via the instability score and yields the predictions used to update them, ensuring we identify and resolve structural inconsistencies in a single auxiliary step.
Formally, we define a score for each token position $i$ that captures its potential to induce large change in the model’s output under perturbations:
$u_{i}=-\,\mathrm{margin}_{i}.$
We then propose the following efficient approximate optimal solution of $\pi$ over a candidate set $C_{t}$:
@@
0: Base model $p_{\theta}$, vocabulary $\mathcal{V}$, prompt $x$, steps NN, step window $\gamma_{s},\gamma_{e})$ where $0\!\leq\!\gamma_{s}\!<\!\gamma_{e}\!\leq\!1$, run revision every $E$ steps, candidate index size $m$, integer remasking limit $k_{\mathrm{rm}}\in\mathbb{N}$ where $k_{\mathrm{rm}}\leq m$
@@
9:   Compute instability: $\ell_{i}\leftarrow-\log p_{\theta}\!\left(y^{(t)}_{i}\mid\tilde{y}^{(t)}\right)$ for all $i\in S_{t}$ {one additional forward pass}
@@
11:   For each $i\in\mathcal{I}_{t}$, set $y^{(t)}_{i}\leftarrow\arg\max_{v}p_{\theta}(v\mid\tilde{y}^{(t)})$ where $v\in\mathcal{V}$
@@ 5.3 Qualitative Results
Figure 4 illustrates how CoRe resolves structural inconsistencies that are otherwise locked in by standard decoding. In this example, the model attempts to initialize a list but commits to a syntax error (“result = = []”). The base LLaDA sampler lacks a mechanism to revisit this decision, effectively anchoring the remainder of the generation to a broken context. In contrast, CoRe detects the high instability of the redundant “=” and the incompatible brackets as the context evolves. It prioritizes these positions for revision, first converting the extra “=” operator to “list” and subsequently adjusting the “[]” brackets to “()”. This process results in the valid expression “list()”, demonstrating CoRe’s ability to break context rigidity and retroactively resolve early structural failures.
@@
Figure 4: CoRe Resolves Structural Inconsistencies Locked by Standard Decoding. The base model commits to a syntax error (“= =”) early in the generation. CoRe identifies the conflicting tokens as context-brittle and invokes revision, successfully recovering the valid contextually stable syntax list().
@@ Appendix C Theoretical Consistency: Computed Instability Lower-Bounds Worst-Case Risk
Fix a step $t$ and let $C_{t}$ be the set of eligible unmasked non-prompt indices. For any subset $S\subseteq C_{t}$, let $\tilde{y}^{(t)}(S)$ denote the perturbed context obtained by masking indices in $S$ (Eq. (1)), and define the instability of index $j$ under this perturbation as
$\ell_{t}(j;S)\triangleq-\log p_{\theta}\!\left(y^{(t)}_{j}\mid\tilde{y}^{(t)}(S)\right).$

### v3

@@ 1 Introduction
Masked Diffusion Models (MDMs) have emerged as a promising alternative to autoregressive decoding for discrete sequence modeling [30], allowing parallel token updates through iterative unmasking [27]. At inference time, an MDM starts with a fully masked sequence and progressively constructs a response. At each diffusion step, the sampler predicts token distributions for masked positions conditioned on the current partially unmasked sequence, selects a subset of positions to unmask, and unmasks them. This iterative process means that unmasked tokens in the early steps are selected under a partial context that evolves as decoding proceeds. To improve quality, some variants employ revision strategies, remasking a small subset of previously unmasked tokens to allow resampling [38].
@@ 2 Related Work
Crucially, these revision methods either rely on unreliable proxy quality signals or require additional training. ReMDM’s standard strategy uses stale confidence scores from past decoding steps, ignoring updated context. P2-Self uses current token probabilities, which can be misleadingly high even for coherent but incorrect generations. Similarly, recent concurrent works like Deferred Commitment Decoding (DCD) [32], Coherent Contextual Decoding (CCD) [6], or Information-Gain (IG) Sampling [39] attempt to mitigate over-confidence by optimizing unmasking order or dynamically delaying commitment. While IG Sampling attempts to minimize future uncertainty by maximizing information gain during selection, it remains a forward-only heuristic. These strategies remain fundamentally passive: they rely on unperturbed internal states and lack a mechanism to revisit previously unmasked tokens, allowing confidently wrong decisions to easily bypass their filters and propagate through the sequence. Moreover, static confidence scores are often uncalibrated [16, 20]. Alternatively, approaches such as PRISM [17], RemeDi [15], GIDD [vonrütte2025gidd] or recent unmasking planners [2] learn specialized policies to guide token revision. Steering or reward-guided methods [29, 37, 26] optimize auxiliary objectives to guide sampling. Although effective, these methods increasingly rely on heavy test-time scaling paradigms or supervised training of external modules. They incur significant overhead through exhaustive search, ensembling, or particle resampling [22, 33], limiting their use as lightweight plug-and-play modules. In contrast, CoRe is context-aware and training-free. Rather than trusting static or stale token uncertainty metrics or passively delaying commitment, we actively probe token stability. We identify context-brittle tokens—those whose likelihood is not stable under context perturbations—framing revision as a distributionally robust optimization problem.
@@ 3 Problem Formulation
Let $\mathcal{V}$ be the vocabulary. Given a prompt $x$, the intermediate MDM sequence at step $t$ is represented by $y^{(t)}=[x;\ y^{(t)}_{1},\ldots,y^{(t)}_{L}\,$], where each $y^{(t)}_{i}\!\in\!\mathcal{V}\cup\{\texttt{[MASK]}\}$. Decoding initializes all $L$ response positions to [MASK]and progressively replaces them with discrete tokens over steps $t=1,\ldots,T$. At each step, for every masked position $i$, the model computes a discrete probability distribution over tokens in the vocabulary:
$p_{\theta}(y_{i}^{(t)}=v\mid y^{(t)},i),\qquad\forall v\in\mathcal{V},$
and the sampler determines which positions to unmask. Unmasking yields the next state $y^{(t+1)}$.
@@ 4 Method
Figure 1: Illustration of Context-Robust Remasking (CoRe). Our method operates on the current state $y^{(t)}$, where the response is partially unmasked. (Left) Candidate Selection: We select potentially brittle unmasked tokens (red dashed box) to test for stability, distinct from the next token scheduled for unmasking (blue dashed box). (Top-Right) CoRe Mechanism: We mask the selected tokens to create a Perturbed Context $\tilde{y}^{(t)}$ and then compute their Instability Scores under this new context. The token “a” is found to be the most brittle (high instability) and is updated to “an,” which is the most likely token given the perturbed context. (Bottom-Right) Base Unmasking: In parallel, the base model uses the original context $y^{(t)}$ to predict the next token (“icy”); this newly unmasked token (“icy”) is combined with the updated token (“an”) to form the Next State $y^{(t+1)}$, yielding the contextually consistent phrase “an icy.”
@@ 4.1 Context-Robust Token Remasking Framework
For a step $t$, let $C_{t}$ be the set of unmasked indices eligible for revision. Let $S_{t}\subseteq C_{t}$ be the indices selected for perturbation (using a selection strategy detailed in Section 4.2). We define the perturbed context $\tilde{y}^{(t)}$ by masking the token positions $i$ in $S_{t}$:
@@
This operation occludes the information at indices $S_{t}$, requiring the model to rely on the remaining context. In our framework, we formally define these masking operations as perturbations to the information content, allowing us to cast revision as a robustness optimization problem.
@@
Let $Y_{i}$ denote the discrete random variable representing the token at position $i$. For each selected position $i\in S_{t}$, we define the instability score $\ell_{i}$ as the negative log-likelihood of the currently generated token $y^{(t)}_{i}$ under the perturbed context $\tilde{y}^{(t)}$:
$\ell_{i}\;\triangleq\;-\log p_{\theta}\!\left(Y_{i}=y^{(t)}_{i}\mid\tilde{y}^{(t)}\right).$ (2)
A high instability score indicates that $y_{i}^{(t)}$ is context-brittle: although the token was originally generated with high confidence, it is inconsistent with the perturbed context $\tilde{y}^{(t)}$. In practice, we use $\ell_{i}$ to rank candidate positions $i\in S_{t}$: positions with larger $\ell_{i}$ are more sensitive to context perturbations (i.e., masking $S_{t}$) and are prioritized for revision. The average instability score over the candidate positions is then:
@@
Having defined the instability loss $\mathcal{L}(S_{t})$ for a fixed subset of tokens, we now formalize the optimization over all possible masking configurations. The masking operation in Eq. (1) produces a family of perturbed contexts by dropping specific token positions, which correspond to the subset $S_{t}$. To formalize the search over these contexts, we introduce a binary mask variable $z\in\{0,1\}^{L}$, where setting $z_{i}=1$ corresponds to masking token $i$; for non-revisable indices $i\notin C_{t}$ we strictly fix $z_{i}=0$. We define the distribution $p_{\pi}(\cdot)$ over these masks as independent Bernoulli trials. The probability of mask $z$ is:
@@
where $\pi_{i}$ controls the probability of perturbing position $i$, and $m$ bounds the expected number of perturbed tokens. To integrate these discrete token representations into our continuous probabilistic framework, we represent the deterministic state of any fixed token $a\in\mathcal{V}$ using the Dirac delta distribution $\delta_{a}(\cdot)$:
@@
This defines a probability distribution that assigns unit mass to the single token $a$, meaning sampling from $\delta_{a}$ always returns $a$ with probability one. Defining the input this way allows the binary mask $z$ to systematically alternate between this known deterministic token and the masked state.
@@
Given a mask vector $z\in\{0,1\}^{L}$, the perturbed sequence $\tilde{y}$ is constructed by conditionally substituting tokens at each position. To facilitate our probabilistic framework, we formalize this deterministic mapping as a degenerate distribution conditional on $z$:
@@
Here, the binary variable $z_{i}$ acts as a routing switch. When $z_{i}=0$, the Dirac delta $\delta_{y^{(t)}_{i}}$ concentrates all probability mass on the original token $y_{i}^{(t)}$; when $z_{i}=1$, the term $\delta_{\texttt{[MASK]}}$ deterministically replaces it with the mask token. Although applying a fixed mask $z$ is a strictly deterministic operation—simply overriding specific tokens with [MASK]—we formulate it as a conditional distribution $Q(\tilde{y}\mid z)$ to allow marginalization over all possible mask configurations. This probabilistic view seamlessly bridges the theory with our algorithm, mapping the binary vector $z$ to the specific set of indices targeted for context perturbation: $S_{t}=\{i\in C_{t}\mid z_{i}=1\}$. The total distribution of perturbed sequences is then the expectation of $Q(\tilde{y}\mid z)$ over all mask configurations:
@@
Step 1: Worst-Case Context Perturbation. We find the worst-case context perturbation by identifying the optimal masking probability vector $\pi^{*}$ that maximizes the expected instability over $S_{t}$:
$\pi^{*}=\operatorname*{arg\,max}_{\pi\in\Pi}\;\Big[\mathbb{E}_{z\sim p_{\pi}(z)}\mathcal{L}(S_{t})\Big],\quad\text{s.t.}\;\;\Pi=\Big\{\pi\in[0,1]^{|C_{t}|}:\sum_{i\in C_{t}}\pi_{i}\leq m\Big\}.$ (8)
Since exactly computing this expectation over the exponentially large discrete space $z\in\{0,1\}^{L}$ is combinatorially intractable, we rely on the tractable approximation detailed in Section 4.2 to identify a certified lower-bound subset.
Step 2: Revision Operation. Given $\pi^{*}$ and a remasking limit $k_{\text{rm}}$, we form the revision set $\mathcal{I}_{t}$ by selecting the $k_{\text{rm}}$ tokens with the highest instability scores under the worst-case context $\tilde{y}\sim Q_{\pi^{*}}$. Computing these scores $\ell_{i}$ yields a certified lower bound on the worst-case instability (Appendix C). Formally, we select:
@@
$y^{(t)}_{i}\leftarrow\operatorname*{arg\,min}_{v\in\mathcal{V}}\Big[-\log p_{\theta}\!\left(Y_{i}=v\mid\tilde{y}^{(t)}\right)\Big],\quad\forall i\in\mathcal{I}_{t}.$ (10)
@@ 4.2 An Efficient Remasking Algorithm
Directly optimizing the worst-case context perturbation in Eq. (8) requires searching over all subsets of token indices and is combinatorially intractable. We therefore consider a single deterministic mask—i.e., a candidate index set $S_{t}$ to perturb. We show in Appendix C that the instability computed on this subset $S_{t}$ constitutes a valid lower bound on the worst-case instability objective. To make this choice tractable, we construct $S_{t}$ based on token contention, specifically targeting positions with the smallest top-2 probability margins:
@@
where $v_{i}^{1}$ and $v_{i}^{2}$ denote the tokens with the highest and second-highest probabilities. Prior work indicates that small top-2 margins reliably signal token contention: competing token values are close in probability [19], meaning even slight context shifts can flip the top-1 prediction which makes them strong candidates to test for context-brittleness.
To formally rank the set of unmasked token positions $C_{t}$ for perturbation, we define the proxy vulnerability score $u_{i}$ as the negative margin:
$u_{i}=-\,\mathrm{margin}_{t}(i).$
We emphasize that while this margin-based proxy identifies where to probe, the revision itself is determined solely by the instability score computed under the perturbed context. This perturbation serves a dual purpose: it quantifies actual brittleness and yields the robust predictions used for updates, ensuring we identify and resolve structural inconsistencies in a single auxiliary step.
To efficiently allocate the expected masking budget $m$ across the set of currently unmasked token positions $C_{t}$, we use a temperature-scaled Softmax to convert these scores into the probability $\pi_{i}$ that each token will be masked. We clip the final value at $1$ to ensure it remains a valid probability:
@@
0: Base model $p_{\theta}$, vocabulary $\mathcal{V}$, prompt $x$, steps NN, step window $\gamma_{s},\gamma_{e})$ where $0\!\leq\!\gamma_{s}\!<\!\gamma_{e}\!\leq\!1$, run revision every $E$ steps, candidate index size $m$, integer remasking limit $k_{\mathrm{rm}}\in\mathbb{N}$ where $k_{\mathrm{rm}}\leq m$; random variable $Y_{i}$ for token at position $i$
@@
9:   Compute instability: $\ell_{i}\leftarrow-\log p_{\theta}\!\left(Y_{i}=y^{(t)}_{i}\mid\tilde{y}^{(t)}\right)$ for all $i\in S_{t}$ {one additional forward pass}
@@
11:   For each $i\in\mathcal{I}_{t}$, set $y^{(t)}_{i}\leftarrow\arg\max_{v}p_{\theta}(Y_{i}=v\mid\tilde{y}^{(t)})$ where $v\in\mathcal{V}$
@@ 5.3 Qualitative Results
Figure 4 illustrates how CoRe resolves structural inconsistencies that are otherwise locked in by standard decoding. In this example, the model attempts to initialize a list but token predictions lead to a syntax error (result == []). The base LLaDA sampler lacks a mechanism to revisit this decision, effectively anchoring the remainder of the generation to a broken context. In contrast, CoRe detects the high instability of the redundant = and the incompatible brackets as the context evolves. It prioritizes these positions for revision, first converting the extra = operator to list and subsequently adjusting the [] brackets to (). This process results in the valid expression list(), demonstrating CoRe’s ability to break context rigidity and retroactively resolve early structural failures.
@@
Figure 4: CoRe Resolves Structural Inconsistencies Locked by Standard Decoding. The base model’s predictions lead to a syntax error (= =) at intermediate decoding step 21. CoRe identifies the conflicting tokens as context-brittle and invokes revision, successfully recovering the valid contextually stable syntax list().
@@ Appendix A Notation
$Y_{i}$ Discrete random variable representing token at position $i$
@@ Appendix C Theoretical Consistency: Computed Instability Lower-Bounds Worst-Case Risk
Fix a step $t$ and let $C_{t}$ be the set of eligible unmasked non-prompt indices. Let $Y_{j}$ denote the discrete random variable representing the token at position $j$. For any subset $S\subseteq C_{t}$, let $\tilde{y}^{(t)}(S)$ denote the perturbed context obtained by masking indices in $S$ (Eq. (1)), and define the instability of index $j$ under this perturbation as
$\ell_{t}(j;S)\triangleq-\log p_{\theta}\!\left(Y_{j}=y^{(t)}_{j}\mid\tilde{y}^{(t)}(S)\right).$